Menu
Waste Processing

Latest news from Calvert Landfill site

March 2015

  • The planting of the 2000 trees on pit 4 have now been completed
  • Site continues to accept spoil trains, up to 3 per day
  • Spoil is being used to continue to restore the site in accordance with the phasing plan
  • Capping is due to start on the latest filled area next month
  • Drilling is also due to start next month

 


 

February 2015

  • Capping of the site for this current phase has now been completed
  • All of the affected gas wells are also reconnected back into the system
  • Site will now have a programme of tree planting starting mid February where up to 2000 trees will be planted on pit4
  • Site continues to fill phase 4 of the over tip which is slower than expected due to the diversions of waste away from landfills
  • Some odour complaints have been received
  • Active area on site is being kept to a minimum and covered at the end of every working day

 


 

January 2015

  • The capping is being completed on site with the last 15,000m2 being placed
  • Due to high winds the site has had to close to some customers to prevent a high environmental impact with regard to litter
  • Both waste and up to 3 spoil trains continue to arrive on site on a daily basis
  • Network rail continue to undertake out of hours maintenance on the rail sidings

 


 

December 2014

  • The final 15m2 of capping has been postponed due to the weather conditions.
  • The works will recommence early 2015, weather permitting
  • Spoil trains continue to arrive on site however there will be a break during the Christmas period with recommencement on the 6th January
  • Waste trains however will continue on a daily basis with the exception of the Holidays
  • Network Rail have informed site that they anticipate undertaking some night time maintenance works on the main line after the Christmas break for a period of time.
  • Those directly affected will have further information posted to them
  • The road repairs to Brackley Lane are likely to commence on the 16th and 17th of December, with Brackley Lane closed form 8am until 4:30 pm both days.
  • Odour complaints continue to reduce with approximately 72% reduction to that of 2013

 


 

November 2014

  • Capping continues on site although weather conditions is making it difficult so the programme has been delayed
  • Tipping in pit 4 phase 4 continues
  • Spoil trains continue to be off loaded at Calvert with an average of 8 per week
  • Waste trains also continue 3-4 per week
  • All additional gas well drilling has ceased due to poor weather making it difficult and unsafe to drill in the identified areas
  • It is expected that the drilling will recommence in the spring of 2015

 


 

October 2014

  • Capping works have been increased by another 30m2 to incorporate parts of pit 4 phase 4 over tip
  • These works will run after the original 50m2 has been completed and will consist of the same material i.e. plastic with 1 metre of spoil on top
  • Tipping continues to be in the rest of phase 4
  • Additional waste trains have been accepted as a result of logistical problems at a waste processing facility
  • As a result the infilling of the next section of the landfill is likely to start at the beginning of 2015
  • Signage to direct traffic to the Brackley Lane entrance is being removed by Transport for Bucks
  • It has been identified that there are 6 signs that need either to be removed or replaced
  • It is hoped that this work can be programmed in before the end of this year

 


 

September 2014

  • The new access road to and from the A41 which is likely to be named Greatmoor Road is now fully open
  • All HGV traffic as form the beginning of the month is now using this road
  • Brackley Lane is no closed to all except Cars, light vehicles and emergency services entering the site
  • Security has been posted at the gate to prevent unauthorised access
  • Additional 24 gas wells are to be installed on site

 


 

August 2014

  • Capping has commenced at the end of August
  • 50m2 of capping is being placed on Phase 3 of the over tip
  • Spoil continues to be placed on the capped areas on phases 1 and 2 ready for final restoration
  • Odour complaints continue to be low
  • Upgrading of gas ring main around site commented, expected to be completed next month
  • Investigation into the works required to allow full adoption by BCC has commenced.
  • Works are unlikely to commence prior to Brackley Lane closure to HGV traffic

 


 

July 2014

  • Site continues to fill pit 4 phase 4 of the over tip
  • Drilling on site continues with most of the 50+ gas wells installed
  • Capping is expected to start in August, slightly delayed
  • Capping will consist of plastic material covered with 1 metre of spoil
  • Spoil trains continue to be delivered to site 2-3 times per day
  • Waste trains continue to be delivered at a rate of 3-4 trains per week

 


 

June 2014

  • Upgrading of the pipeline into the gas utilisation plant is now completed along with a section that was removed, cleaned and replaced
  • During this time additional flares were installed to ensure adequate gas extraction
  • Some complaints have been received with regard to speeding traffic coming to the site. Although FCC have no legal authority in ensuring customers to the facility adhere to the highway code, if any information is given to site we will inform our customers of the concerns raised
  • Odour complaints continue to be low

 


 

May 2014

  • A programme to install 55 gas extraction wells is currently being prepared with installation expected to start next month
  • These wells will be installed in areas which will be capped.
  • Other wells will be drilled to replace some that have been decommissioned
  • Decommissioning of the wells identified is continuing
  • Expected capping works to start in June or July

 


 

April 2014

  • Continued reduction of odour complaints received by site and EA
  • Additional spoil being placed on phases 1 and 2 of the over tip ready for tree planting expected to start late 2014
  • Waste continues to be places in phase 4 of the over tip
  • Void calculation suggest that this area will continue to be filled until March 2015
  • Camera survey carried out early in the year identified some wells that are no longer viable
  • A programme of works is being prepared to decommission these wells over the next few months

March 201

  • Pit 4 phase 4 over tip continues to be filled
  • Site continues to accept spoil trains
  • The stockpile for the engineered bund continues to be added to from the spoil arriving on the trains
  • Reduction in waste trains to site due to diversions away from landfill in operation
  • Engineering works expected to start later in this year which will be the capping of phase 3 and parts of phase 4 of the over tip

February 2014

 
  • The rain continues
  • The road sweeper continues to operate on, around and off site
  • A specialised road sweeper continues to clean the external road on a Saturday
  • Wheel wash surveys continue to ensure all HGVs leaving site are using it correctly
  • Some odour complaints received, a similar number to January
  • Due to the poor weather conditions works to the gas infrastructure has been difficult
  • However gas wells continue to be monitored

January 2014
  • Continued rain making the site very difficult to work
  • Additional road sweeper employed on Saturday to aid in the cleaning of the external road
  • Temporary repairs undertaken in Brackley Lane due to the presence of pot holes
  • Waste train deliveries in the morning have now stopped
  • Spoil trains continue to be accepted
  • Odour complaint continue to decrease compared with 2013

December 2013
  • Increase complaints with regard to mud on the external road.
  • The road sweeper is continuing to sweep the road to ensure that it is maintained
  • The EA have reviewed the site activites and are happy with all site actions to prevent any debris on the roads
  • A wheel wash survey has taken place to review the users of the wheel wash to ensure that they are using it correctly
  • Result show that on the several occasions all were observed using it correctly
  • Observations were from a discreet distance
  • Odour complaints have decreased

November 2013
  • The rain continues to fall resulting in increase difficulties on site
  • The road sweeper has been busy sweeping the internal and external roads to aid in the prevention of debris being tracked off site
  • Some odour complaints received as a result of the wet waste arriving in the morning on the train
  • The deodorising system continues to be fully functional
  • Improvements have been made to the gas pipe work infrastructure to aid in the odour control on site

October 2013
  • High levels of rain fall on site has resulted in operational difficulties
  • As a result the tipping areas are kept to a minimum and the dipping distances have been kept short to allow continued access to the active area
  • Odour complaints have decreased since last month
  • Site continues to accept waste and spoil trains
  • The stockpile for the engineered bund continues to be added to from the spoil arriving on the trains

September 2013
  • Additional leachate wells drilled on site, resulting in localised odours. Off site was checked for odours associated with the landfill and this activity along with an FID survey. Nothing was found
  • The EA also undertook an offsite olfactory survey during this time and confirmed no odours were detected as a result of the land filling operations
  • The drilling and connection of the leachate wells were complete at the end of the month
  • Agriculture odours have also continued to be detected off site during the month of September
  • Results of a site FID survey carried out this month indicates that the capping undertaken at the beginning of this year continues to ensure a high level of environmental compliance by preventing elevated readings

August 2013
  • Additional leachate pumps installed on site
  • Horizontals continue to be installed within the active areas
  • Due to difficulties with a train haulier, there was a 24 hour delay on off loading a waste train. As a result the waste, when disposed of on site was stronger in odour than normal
  • The deodoriser was, as a result, operational constantly whist the train was being off loaded and covered with suitable inter material

July 2013
  • Due to high temperatures experienced the train waste has been identified as being stronger in odour than normal as a result of the waste ‘cooking’ in the containers
  • As a result the train waste was deposited on site without delay with daily covered applied once complete
  • The tipping area is kept to a minimum operational size also to assist in the reduction of odours associated with this waste
  • The additional 25 gas wells have been completed all are now connected into the gas ring main
  • The upgrade of the gas ring main has been completed
  • Horizontal pipe work continues to be installed within the active area to assist in the environmental controls on site
  • Towards the end of the month significant external odour sources have been identified confirmed independently by the Environment Agency

June 2013
  • The drilling of the 44 gas and leachate wells has been completed
  • All of these wells are now connected into the gas collection system
  • The upgrading of the ring main has been completed
  • Additional 25 gas wells are being drilled in the recently finished capped area on pit 4
  • These are being connected into the gas collection system without delay
  • Site continues to import soil on the train for the stability bund requirements in pit 6
  • Additional evening trains expected to arrive on site at 17:00 and depart at 23:00
  • Off loading procedure adapted for this evening train to assist in the reduction of noise noted from the locomotive
  • FID survey of the site carried out to assist in the identification of any potential gas emissions from site
  • Action plan in place to address any issues
 

May 2013
  • All capping has been completed on site, including the area that has been brought forward from 2014
  • All cap stripping is now complete for 2013 providing enough void space until mid 2014
  • Drilling of 44 gas and leachate wells has commenced with 3 drilling rigs on site to assist in the programme
  • The wells are connected without delay into the gas extraction system as they are installed
  • The wells are situated in 3 main locations; on the flank of pit 5, beside the main haul road and where the capping has just been completed
  • This will bring the total number of gas wells on site to in the region of 650
  • Upgrading of the gas ring main has commenced
  • Some odours noticed on site were as a result of offsite agricultural activities

April 2013
Update on Capping:
  • The commencement of capping on phase 4 has been pushed forward.
  • In accordance with the 5 year phasing plan it was scheduled for 2014, however due to slight changes in priority some of phase 4 is now filled to height and is being capped.
  • The regulating layer is in place on the flank of phase 3 and parts of phase 4 of the over tip on Dunsty Hill The first layer of the permanent capping is in place on phase 3. The layer is now been checked for permeability in accordance with the CQA plan before continuing with the next layers.
  • The access road along the flank of phase 3 and 4 is being stripped out and capped without delay. Works on this is expected ot be complete by end April.
  • Temporary capping is also ongoing on other areas of Pit 4.
  • A secondary FID survey has been undertaken to compare with the survey carried out by the EA last month. This allows the remediation actions to be checked.
  • During the capping process gas wells are disconnected, and reconnected without delay after the works around that well are complete.
  • This was the case at the beginning of April; the wells were disconnected in the morning and were reconnected by 15:00 the same day.
  • Due to the recent high winds, litter has collected on site around the edge of the current tipping area. This is being picked on a daily basis as required
  • Weekend works to repair the internal haul road around the weighbridge was undertaken. These works are now complete
     

     
March 2013 
  • The current tipping area continues to be covered with suitable inert material at the end of every working day.
  • Horizontals installed and connected in active areas.
  • Contractors on site (18th) have started the capping works on Pit 4 phase 3.
  • They are prioritising areas with high odour potential.
  • They are working subject to weather conditions Monday to Friday 12 hours a day and Saturday until 4.00 pm in order to do as much capping as possible.
  • Gas well disconnection and connection will be carried out as efficiently as possible to minimise any potential impact.
  • The contaminated surface water pool has been emptied.  When the weather permits it will be completely filled with clay
  • Following the FID survey on site a list of infrastructures has been identified and an action plan will be put together to rectify and replace any installations if required.
 

 
February 2013
  • The current tipping area continues to be covered with suitable inert material at the end of every working day.
  • The deodorising system is fully operational although it has been affected during by the low temperatures. It was also noted that on one Saturday evening the system had tripped. It was restarted and a repeat of this had not occurred.
  • Horizontals installed and connected in active areas.
  • Site detected waste odours off site during the daily Odour Assessment between 8 and 8:10am, rechecked approximately 2 hours later, no odours detected.
  • A FID was carried out at the time of the detection there were no readings off, or on, the site boundary.
  • Intensive FID survey carried out around site, and ongoing, and findings reported.  Remediation works ongoing as appropriate.
  • Works to pump some contaminated water from a pool in pit 4 ongoing, once dry it will be sealed with clay.
 

 
January 2013
  • The current tipping area has been relocated to assist in reducing the likelihood of odours.
  • The internal haul road to the tipping face has been relocated and old road filled over with appropriate horizontal installed.
  • Horizontals installed and connected in active areas.
  • Gas wells are connected as quickly as possible, and wells which are known to be causing an environmental impact are temporarily connected to the gas system at the end of every day and disconnected during operational hours to allow access to tipping area.
  • Due to freezing conditions the deodorising system has not been operational.  However, will be fully operational as soon as weather conditions allow.
x Staff zone login

Please enter your details below

 
 
 
Submit